OpenStudy (anonymous):

Mona has a triangular flag. She wants to cut out a similar flag that has dimensions one-fourth of the original flag. If the area of the original flag is 320 square inches, what is the area of the new flag that Mona cuts out? a.16 in2 b.20 in2 c.32 in2 d.80 in2

OpenStudy (anonymous):

if each dimension is reduced by 1/4th the ur new base and ht will be b/4 and h/4. so the new area is = 0.5 * b/4 * h/4 = 1/16 ( 0.5 * b * h) but u already know that 0.5*b*h = 320. so new area = 320/16 = 20.

OpenStudy (allank):

As Vaidehi09 said, Calculate the linear scale factor (LSF) "dimensions one-fourth of the original flag" LSF = 1/4 Calculate the area scale factor (ASF) ASF = LSF^2 = (1/4)^2 = 1/16 The area of the new flag will be 1/16 of the original.

OpenStudy (anonymous):

Okay i see ! thankyouu @Vaidehi09 @allank

OpenStudy (allank):

You're welcome :)

OpenStudy (anonymous):

sure, np!